Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
yutaro.enomoto - 21:29 Tuesday 18 December 2018 (7425) Print this report
ALS noise investigation: ADC noise and re-measurement of VCO noise

Today I wanted to re-measure the VCO noise since we have changed the frequency reference for the low noise VCO,
and during this measurement process found that measured spectra of out-of-loop stability of ALS (7256, for example) are limited by ADC noise below 10 Hz.

= ADC noise =
The first figure shows the control signal (SLOW_DAQ or SLOW_MON) of green lock when the main laser was locked to the arm cavity, which corresponds to out-of-loop noise of ALS, and ADC noise plotted with them.
This clearly shows that low frequency below 10 Hz was limited by ADC noise.
We have to do the similar measurement again with sufficient amplification applied to the control signal to monitor it with better SNR.

= VCO noise measurement again =
Here is the measured noise we got today, shown along with the previous result.
As can be seen, today's result shows the lower noise. I don't know why.
The differences between the previous setup and today's setup are:
-- Operating point of VCO
Previous: 1.5 V
Today: 1.75 V (close to the operating point of VCO when ALS CARM is closed)
-- Frequency reference for VCO
Previous: Agilent N5181B
Today:Keysight 33611A, option OCX
Both should be stable enough and should not limit the VCO noise.

Images attached to this report
Comments to this report:
yutaro.enomoto - 10:47 Wednesday 19 December 2018 (7432) Print this report
Previous measurement 6552
Search Help